Find a particular solution of the differential equation: ((2y-x))/((y+2x) ) ⅆy/ⅆx=1 given that y=3 when x=2


1
Expert's answer
2022-03-08T13:09:01-0500

"\frac{2y-x}{y+2x}\frac{dy}{dx}=1, y=3, x=2.\\
\frac{dy}{dx}=\frac{y+2x}{2y-x}\\
y=xz(x), y'=xz'+z\\
xz'+z=\frac{xz+2x}{2xz-x}\\
xz'+z=\frac{z+2}{2z-1}\\
xz'=\frac{z+2}{2z-1}-z\\
xz'=\frac{-2z^2+2z+2}{2z-1}\\
\frac{2z-1}{-2z^2+2z+2}dz=\frac{dx}{x}\\
-\frac{z-\frac{1}{2}}{z^2-z-1}dz=\frac{dx}{x}\\
-\frac{z-\frac{1}{2}}{z^2-2z\frac{1}{2}+\frac{1}{4}-\frac{1}{4}-1}dz=\frac{dx}{x}\\
-\frac{z-\frac{1}{2}}{(z-\frac{1}{2})^2 -\frac{5}{4}}dz=\frac{dx}{x}\\
-\frac{d((z-\frac{1}{2})^2 -\frac{5}{4})}{(z-\frac{1}{2})^2 -\frac{5}{4}}=2\frac{dx}{x}\\
-\int\frac{d((z-\frac{1}{2})^2 -\frac{5}{4})}{(z-\frac{1}{2})^2 -\frac{5}{4}}=2\int\frac{dx}{x}\\
-\ln|(z-\frac{1}{2})^2 -\frac{5}{4}|=2\ln|x|+\ln|C|\\
\frac{1}{(z-\frac{1}{2})^2 -\frac{5}{4}}=Cx^2\\
\frac{1}{z^2-z-1}=Cx^2\\
\frac{1}{(\frac{y}{x})^2-\frac{y}{x}-1}=Cx^2\\
\frac{1}{(\frac{3}{2})^2-\frac{3}{2}-1}=4C\\
C=-1\\
\frac{1}{(\frac{y}{x})^2-\frac{y}{x}-1}=-x^2\\"